OpenStudy (anonymous):

Will give medal. c: Solve for x: |2x + 6| − 4 = 20

OpenStudy (anonymous):

start by isolating the absolute value

OpenStudy (aum):

|2x + 6| − 4 = 20 To solve for x, isolate x. Add 4 to both sides. What do you get?

OpenStudy (anonymous):

|2x+6| - 4= 20 +4 +4 |2x+6|= 24 ?

OpenStudy (anonymous):

once you've done that make your two different problems and solve for x for each. Your first one will be 2x + 6 = 24 and your second will be 2x + 6 = -24

OpenStudy (aum):

Yes. Use this fact: If |A| = B, it means A = B, A = -B

OpenStudy (anonymous):

okay I'll write that down

OpenStudy (anonymous):

once you solve for x for each, that will give you your x values

OpenStudy (aum):

If |A| = B, it means A = B, A = -B Therefore, |2x+6|= 24 means 2x+6 = 24 or 2x+6 = -24 Solve each equation separately

OpenStudy (anonymous):

2x+6= 24 -6 -6 2x= 18 2x/2 = 18/2 x= 9 2x+6= -24 -6 -6 2x= -30 2x/2= -30/2 x= -15 did i do this right?

OpenStudy (aum):

Yes!

OpenStudy (anonymous):

yay thanks!

OpenStudy (aum):

You are welcome.